OpenStudy (anonymous):

Plz help mid life problem here!!!! Find the standard form of the equation of the parabola with the given characteristic(s) and the vertex at the origin.....??????

OpenStudy (anonymous):

y = x^2

OpenStudy (anonymous):

How can u plZ explain.?????

OpenStudy (anonymous):

If you plot y against x for y = x^2 When x = 0, y = 0 When x = 1, y = 1 When x = 2, y = 4 When x = -1, y = 1 When x = -2, y = 4 Plot these points on a graph, and you will obtain the required parabola.

OpenStudy (anonymous):

Wait the directrix is y=3

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

The vertex of a parabola of the form \(y=ax^2+bx+c\) is located at the following point.\[\huge \text{Vertex}=\left(-\frac{b}{2a},\ c-\frac{b^2}{4a}\right)\]So, come up with \(a\), \(b\), and \(c\) that gives a vertex \((0,0)\). (really easy)

OpenStudy (anonymous):

I honestly do not understand

OpenStudy (anonymous):

Sorr:(

OpenStudy (anonymous):

Y

OpenStudy (anonymous):

Peas help me!

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Plz,!!!!!!!!